What is the PH of a solution made by mixing 30.00
Answer #1

27.

First we need to obtain the moles for each compound

30 mL\frac{0.10 mol HCl}{1000 mL}=3x10^{-3}moles HCl

40 mL\frac{0.10 mol KOH}{1000 mL}=4x10^{-3}moles KOH

Then we write the process

HCl KOH
Initial 0 4x10-3 mol
Change
3x10-3 mol
---
Eq 0 1x10-3 mol

Then dividing the moles of KOH between the total volume:

\frac{1x10^{-3}}{70 mL}\frac{1000 mL}{L}=0.01429 M KOH

Using the following equation we get:

pOH=-log[OH]

pOH=-log[0.01429 M]

pOH=1.845

Using the relation between pOH and pH

pH+pOH=14

pH=14-1.845

pH=12.15

The answer is A) 12.15

We are only allowed to answer 1 question per Q&A, please consider posting multiple question in multiple sets of Q&A.